OpenStudy (anonymous):

Blake has a total of 64 DVD's and CD's. If the number of CD's is eight more than three times the number of DVD's, how many DVD's does he have?

OpenStudy (anonymous):

is it 14?

OpenStudy (anonymous):

LET: C - NO. OF CD --------D - NO. OF DVD C + D = 64 EQ.1, C = 64 - D C = 3D + 8 EQ. 2 EQUATE EQ.1 AND 2 64 -D = 3D + 8 SOLVING FOR D = ?

OpenStudy (anonymous):

50 cd's and 14 dvd's

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

= 14 NO. OF DVD CORRECT!!

OpenStudy (anonymous):

Awesome, thanks!

OpenStudy (anonymous):

Your welcome. c: